the area of a triangle flag is 30 square centimeters. Its altitude is 2 centimeters longer than twice its base. FInd the lengths of the altitude and the base.
Answer by TimothyLamb(4379)

---
area = (1/2)bh
b = x
h = 2x + 2
---
30 = (1/2)x(2x + 2)
60 = 2xx + 2x
2xx + 2x - 60 = 0
---
the above quadratic equation is in standard form, with a=2, b=2, and c=-60
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
2 2 -60
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
this quadratic has two real roots (the x-intercepts), which are:
x = 5
x = -6
---
the negative root does not fit the problem statement, so use the positive root
---
answer:
b = 5 cm
h = 12 cm
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php